OpenStudy (anonymous):

solve each equation by hand using only algebra 7/3(2x-1)=1/5x+2/5(4-3x)

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{7}{3(2x-1)} =\frac{7}{ \frac{5x+2}{5(4-3x)} } }\) like this ?

OpenStudy (solomonzelman):

I mean \(\LARGE\color{blue}{ \frac{7}{3(2x-1)} =\frac{1}{ \frac{5x+2}{5(4-3x)} } }\)

OpenStudy (anonymous):

@SolomonZelman no the numbers where there is a slash its just a fraction alone the things in parenthesis are right next to them like I have them written above

OpenStudy (solomonzelman):

can you draw the problem then ?

OpenStudy (anonymous):

yes just give me one moment

OpenStudy (solomonzelman):

sure

OpenStudy (anonymous):

@SolomonZelman \[\frac{ 7 }{ 3 }\left( 2x -1 \right)=\frac{ 1 }{ 5 }x +\frac{ 2 }{ 5 }\left( 4-3x \right)\]

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \frac{7}{3}(2x-1) = \frac{1}{5}x+\frac{2}{5}(4-3x) }\) \(\Large\color{blue}{ \frac{7}{3}(2x-1) = \frac{x}{5}+\frac{2(4-3x)}{5} }\) \(\Large\color{blue}{ \frac{7}{3}(2x-1) = \frac{x+2(4-3x)}{5} }\) \(\Large\color{blue}{ \frac{7}{3}(2x-1) = \frac{x+8-6x}{5} }\) \(\Large\color{blue}{ \frac{7}{3}(2x-1) = \frac{8-5x}{5} }\) \(\Large\color{blue}{ \frac{7(2x-1)}{3} = \frac{8-5x}{5} }\) \(\Large\color{blue}{ \frac{14x-7}{3} = \frac{8-5x}{5} }\)

OpenStudy (solomonzelman):

From here, I suggest to cross multiply.

OpenStudy (anonymous):

@SolomonZelman what would be the answer if you crossed multiplied

OpenStudy (solomonzelman):

5(14x-7)=3(8-5x) 70x-35=24-15x 70x+15x=24+35 85x=59

OpenStudy (solomonzelman):

Looks like it is 59/85